OpenStudy (anonymous):

Factor the polynomial. x3 - 2x2 + x - 2 Hint: use " i " to enter factors containing square roots of negative numbers. For example, if your factor is (x + ), you should write (x + 2i).

OpenStudy (anonymous):

$$ x^3-2x^2+x-2=x^2(x-2)+(x-2)=(x-2)(x^2+1)=(x-2)(x+i)(x-i)$$